A hydraulic lift has input and output pistons with diameters of 0.12 m and 0.48 m, respectively. If a force of 275 N is exerted at the input piston, calculate the maximum mass that can be lifted by the output piston.


1
Expert's answer
2022-02-11T08:49:31-0500

Given:

"F_1=275\:\rm N"

"d_1=\rm 0.12\: m"

"d_2=\rm 0.48\: cm"


The hydraulic press equation says


"\frac{F_1}{A_1}=\frac{F_2}{A_2}"

Hence, the maximum mass that can be lifted by the output piston


"m=F_2/g=F_1\frac{A_2}{gA_1}\\
=275*\frac{(0.48)^2}{9.8*(0.12)^2}=449\: \rm kg"